# A Neurosymbolic Framework for Interpretable Cognitive Attack Detection in Augmented Reality

### v1

@@ Abstract
Augmented Reality (AR) enriches perception by overlaying virtual elements on the physical world. Due to its growing popularity, cognitive attacks that alter AR content to manipulate users’ semantic perception have received increasing attention. Existing detection methods often focus on visual changes, which are restricted to pixel- or image-level processing and lack semantic reasoning capabilities, or they rely on pre-trained vision–language models (VLMs), which function as black-box approaches with limited interpretability. In this paper, we present CADAR, a novel neurosymbolic approach for cognitive attack detection in AR. It fuses multimodal vision–language inputs using neural VLMs to obtain a symbolic perception-graph representation, incorporating prior knowledge, salience weighting, and temporal correlations. The model then enables particle-filter based statistical reasoning–a sequential Monte Carlo method–to detect cognitive attacks. Thus, CADAR inherits the adaptability of pre-trained VLM and the interpretability and reasoning rigor of particle filtering. Experiments on an extended AR cognitive attack dataset show accuracy improvements of up to $10.7\%$ over strong baselines on challenging AR attack scenarios, underscoring the promise of neurosymbolic methods for effective and interpretable cognitive attack detection.
@@ 1 Introduction
Augmented Reality (AR) and Mixed Reality (MR) technologies have rapidly advanced in recent years, offering users an immersive perceptual experience by overlaying contextually relevant information and virtual objects onto their physical environment (Zheng and Yuan 2023). As AR/MR are rapidly changing how users perceive, interpret, and interact with the world around them, new vulnerabilities have emerged and attracted increasing research attention. In particular, the reliance on visual and spatial cues in AR/MR creates an implicit trust in the accuracy and authenticity of overlaid content. This trust can be exploited by adversaries to launch cognitive attacks, which use malicious manipulations of virtual content to mislead, distract, or coerce users without their awareness (Xiu, Scargill, and Gorlatova 2025; Cheng et al. 2023; Teymourian et al. 2025). Unlike traditional cybersecurity threats that target data or system integrity, cognitive attacks in AR/MR (such as the examples shown in Fig. 1) operate at the perceptual level, manipulating users’ semantic interpretation through subtle or deceptive visual alterations (Xiu and Gorlatova 2025).
@@
Figure 1: Illustrative AR/MR cognitive attacks. Top row: original scenes; bottom row: manipulated scenes. (a) Text Modification Attack. (b) Visual Modification attack. (c) Removal Attack. (d) Addition Attack.
@@
Prior work on detecting such cognitive attacks often employs computer vision techniques to identify visual changes or adversarial distortions (Qiu et al. 2019; Duan et al. 2022a, b; Li et al. 2024; Acharya et al. 2022). However, these methods rely on pixel/image processing, lacking the ability to model the human perception process that not only involves prior knowledge and contextual information, but also selectively focuses on aspects of the visual field for semantical interpretation. As a result, these methods may struggle to capture smart cognitive attacks manipulating only a small number of pixels and also could suffer from false positives due to benign visual perturbations that do not affect users’ semantic interpretation (Li et al. 2024). More recent studies propose pre-trained vision–language models (VLMs) (Xiu, Scargill, and Gorlatova 2025; Roy et al. 2025) to achieve semantic-level detection, by processing visual features and incorporating human-like knowledge. But the reasoning of pre-trained VLMs is a black-box approach offering limited interpretability. It is hard to leverage domain-specific knowledge or model scene/context evolutions or temporal correlations in cognitive attack detection.
@@
In this paper, we consider cognitive attacks that aim to manipulate users’ semantic perception in AR/MR. We propose CADAR (Cognitive Attack Detection in Augmented Reality), a neuro-symbolic approach encompassing two key components working in concert: (i) a symbolic perception-graph model, which is extracted using pre-trained neural VLM models to represent scene/context evolution at semantic-level, ; and (ii) a particle-filter model, which is commonly used in non-linear and stochastic systems. It applies a sequential Monte Carlo method to the perception graph for reasoning and detecting cognitive attacks, stabilizes inherently noisy or inconsistent VLM outputs, and mitigates errors or missed detections through probabilistic fusion. It also offers granular, flexible temporal detection: by using reference sequences, frame-wise inference localizes attacks per frame, to produce a variable number of labels when needed. Thus, the CADAR approach inherits the adaptability of pre-trained VLM (neural part), as well as the interpretability and reasoning rigor of particle filtering (symbolic part). It semantically interprets and models the sequential scene changes for robust cognitive attacks detection in AR/MR, overcoming limitations in existing work.
Our key contributions are as follows:
@@
  • Neuro-symbolic modeling and reasoning. We present a neuro-symbolic framework tailored to AR/MR cognitive attack detection. A pre-trained neural VLM generates symbolic perception graphs that stores objects, attributes, and relations for explicit, interpretable reasoning.
  • Perception graph representation. We propose a perception graph schema that formally encodes multi-modal scene/context evolution at the perceptual level, integrating outputs from complementary modules (image and text encoders, object detection, and segmentation) for cognitive attack inference.
  • Particle-filter-based statistical reasoning. We are the first to apply particle filtering to symbolic perception graphs. Not only does our method improve the robustness of stochastic VLM outputs, mitigating noise and hallucination, but it also enables interpretable statistical procedures to reason about cognitive attacks.
  • We create the Extended AR-VIM dataset, which is the first publicly available AR cognitive attack dataset, by expanding the existing AR-VIM dataset (Xiu and Gorlatova 2025). CADAR evaluation shows it significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model.
@@
The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 presents our method; Section 4 reports experimental results and Section 5 concludes the paper.
@@ 2 Related Work
One well-studied category of visual attacks is adversarial attacks, which introduce imperceptible yet malicious perturbations to input data to mislead machine learning models (Zhang et al. 2024a; Mumcu and Yilmaz 2024). In contrast, cognitive attacks target perception-level semantics and aim to alter a human user’s understanding of a scene (Cheng et al. 2023; Xiu and Gorlatova 2025). Augmented Reality (AR) systems, which overlay virtual content onto the physical world to enhance perception and decision-making, are particularly vulnerable to such attacks. However, research on cognitive attacks in AR remains limited. Existing methods often rely on access to the original, unaltered scene for comparison or employ pre-trained VLMs with limited interpretability.
Knowledge Graphs.
Graph-based representations offer a compact and expressive way to encode entities and the possibly complex relations among them, making graphs well-suited for downstream prediction and retrieval tasks. A knowledge graph represents each entity as a node and each semantic relation as an edge, thereby capturing structured world knowledge at scale (Hogan et al. 2021). A scene graph extends this idea to visual understanding: it organizes objects, their attributes, and inter-object relationships into a graph that supports higher-level reasoning over images and videos (Chang et al. 2021). Numerous task-specific graph formulations have also emerged, for example, mission-specific knowledge graphs for situational awareness (Yun et al. 2025), topological graphs for robot navigation (Shah et al. 2022), and interaction networks that model physical relations among objects (Battaglia et al. 2016).
@@
Neuro-symbolic approaches couple deep neural networks with symbolic reasoning–i.e., reasoning over explicitly represented knowledge in formal languages, to capitalize on the complementary strengths of subsymbolic pattern recognition and symbolic compositionality (Bhuyan et al. 2024). Within knowledge-graph research, most neuro-symbolic work addresses link prediction or graph completion, where the hybrid models offer interpretable rules, guided learning, and more transparent decision pathways (DeLong, Mir, and Fleuriot 2024). Yet, like standard supervised techniques, many of these solutions still demand sizable labeled datasets. In contrast, our framework introduces a neuro-symbolic reasoning module that combines explicitly encoded rules with statistics, achieving high interpretability while remaining data-efficient.
@@
Particle filters–sequential Monte Carlo algorithms for recursive Bayesian estimation–propagate a set of weighted samples to approximate the posterior state distribution, outperforming classical Kalman filters in strongly non-linear or non-Gaussian settings (Künsch 2013). They are widely used in surveillance, robotics, and navigation, leveraging visual, geometric, and motion cues such as color, texture, and shape (Awal et al. 2023). In CADAR, we model graph nodes and edges as stochastic particles, enabling robust reasoning under noisy or inconsistent VLM outputs.
@@ 3 Methodology
Figure 2: CADAR system overview. Sequential video frames are converted into temporal perception graphs, which the attack-detection module then analyzes to identify, classify, and localize adversarial attacks.
@@ 3.1 Cognitive Attacks
We study four representative adversarial attack types that compromise an AR scene’s semantics (Fig. 1).
Text-modification Attack.
Altering on-scene text, e.g., overlaying “Free Parking” atop a “No Parking” sign, reverses instructions or injects false information, misleading both users and perception systems.
Visual-modification Attack.
Altering an object’s appearance or position (green light to red, stop sign moved to mid-junction) distorts recognition and user judgment.
Removal Attack.
Occluding or deleting safety-critical cues (e.g., an exit sign) erases vital information and breaks expected graph relations.
Addition Attack.
Inserting fictitious objects or labels (e.g., a fake hazard symbol) plants misleading cues that divert attention and corrupt downstream reasoning.
@@ 3.2 Perception Graph Model
Figure 3: Perception Graph generation: Given a video frame and its contextual description, generate the corresponding perception graph at time step t.
@@
A perception graph consists of nodes and edges. Objects are represented as nodes, and object relationships are represented as edges. Given the context information and prompt (shown in the Figure  3), the VLM would extract the related objects with their relationships in the scene. At timestep $t$, the Perception Graph $G_{t}$ is generated as follows.
Graph Construction.
@@
Thus the GraphSet $\hat{G}_{t}$ is the union of all spatial and temporal connections that stitch object instances across time steps 1:t1{:}t. Each node still carries its original timestamp to preserve the temporal ordering.
@@
$\hat{G}_{t}=\bigl{(}\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}\bigr{)},\quad\mathcal{V}_{1:t}=\!\!\bigcup_{i\in\mathcal{I}_{t}}\!\!\mathcal{N}_{i,t},\quad E_{1:t}=\!\!\bigcup_{i,j\in\mathcal{I}_{t}}\!\!\mathcal{E}_{ij,t}.$ (7)
@@ 3.3 Particle Filter Model
Our particle filter model tracks graph evolution and analyzes adversarial perturbations. It proceeds in three stages: prediction and matching, attack detection, and state estimation (cf. Fig. 4). Each entity (node or edge) is treated as a particle At each timestep, particles match the incoming observations with prior particles. The attack detection algorithm applies statistical tests to newly matched particles and discards those marked as attacked. Finally, weighting, resampling, and estimation produces an updated, temporally consistent graph state, preserving persistent NodeSet/EdgeSet identities and the relational structure over time.
@@
$s_{j,i,m}\;=\;\frac{1}{|\mathcal{N}_{i,t-1}|}\sum_{v\in\mathcal{N}_{i,t-1}}\operatorname{sim}_{m}(v_{j}^{t},v),\qquad m\in\{n,d,f\}.$ (9)
@@
For each remaining candidate, measure the distance
$D_{j,i}\;=\;\sum_{m\in\{n,d,f\}}\bigl{|}s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr{|}.$ (11)
Finally, the assignment rule:
$\mu_{t}(v_{j,t})=\begin{cases}\displaystyle\arg\min_{i\in\mathcal{C}_{j}}D_{j,i},&\mathcal{C}_{j}\neq\varnothing,\\[6.0pt] \text{new\_id},&\mathcal{C}_{j}=\varnothing.\end{cases}$ (12)
If $\mu_{t}(v_{j}^{t})$ returns an existing ID, append $v_{j}^{t}$ to that NodeSet and update $A_{i}^{t}$; otherwise spawn a new NodeSet. Confirmed edges $e_{ab}^{t}$ are then inserted into $E_{\mu_{t}(a)\,\mu_{t}(b),t}$ as before.
After matching, the NodeSet goes into the Attack detection process.
Attack Detection.
@@
where $k_{m}$ is chosen such that $\geq\!99\%$ variance is retained. The projected embeddings are
$\mathbf{z}_{m}(v)=\mathbf{W}_{i,m}^{\,t-1\,\top}\mathbf{m}(v)\;\in\mathbb{R}^{k_{m}}.$ (14)
For each modality, fit the Gaussian reference model
$\mathbf{z}_{m}(v)\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{c}_{i,m}^{\,t-1},\,\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr{)},\quad v\in\tilde{\mathcal{N}}_{i,t-1},$ (15)
@@
$\mathbf{c}_{i,m}^{\,t-1}=\frac{1}{|\tilde{\mathcal{N}}_{i,t-1}|}\sum_{v\in\tilde{\mathcal{N}}_{i,t-1}}\mathbf{z}_{m}(v),$ (16)
$\boldsymbol{\Sigma}_{i,m}^{\,t-1}=\mathrm{Cov}\!\bigl{\{}\mathbf{z}_{m}(v)\bigr{\}}.$ (17)
@@
For the structural attacks, specifically object removal (Type 3) and object addition (Type 4), the detector relies on lightweight temporal heuristics rather than statistical distance. After each matching step the algorithm keeps a two-frame sliding window of NodeSet activity. If an object that previously carried a high task-importance score ($\pi\geq\pi_{\text{high}}$) fails to obtain a match in both the current and the immediately preceding frame, the system labels the disappearance a removal attack. Conversely, any NodeSet that first appears within the last two frames and whose reasonability score falls below a predefined threshold ($\rho\leq\rho_{\text{low}}$) is treated as an addition attack.
@@
So $w_{j}$ is the normalized weight of the particle, based on its distance to the new particle. Then use the weight to resample $M$ indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ yields
$\hat{a}\;=\;\frac{1}{M}\sum_{k=1}^{M}a^{(s_{k})},$ (22)
@@
For embedding attributes, modality $m\in\{n,d,f\}$ (name, description, image) let $\mathbf{m}^{\star}=\mathbf{m}(v_{i}^{t})$ and $\mathbf{m}^{(j)}=\mathbf{m}(v^{(j)})$. Cosine similarity gives
$sim_{j}=\frac{\mathbf{m}^{(j)}\!\cdot\!\mathbf{m}^{\star}}{\|\mathbf{m}^{(j)}\|\,\|\mathbf{m}^{\star}\|},\qquad w_{j}=\frac{\exp(\lambda_{m}sim_{j})}{\sum_{k=1}^{M}\exp(\lambda_{m}sim_{k})},$ (23)
with a scale factor $\lambda_{m}>0$. Resampling as above yields indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ and the unit-normalised mean
$\hat{\mathbf{m}}\;=\;\frac{\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}}{\Bigl{\|}\sum_{k=1}^{M}\mathbf{m}^{(s_{k})}\Bigr{\|}},$ (24)
would be the resampled embedding particle.
@@
$\hat{v}_{i,t}=\bigl{(}\hat{p},\;\hat{\pi},\;\hat{\rho},\;\hat{\mathbf{n}},\;\hat{\mathbf{d}},\;\hat{\mathbf{f}}\bigr{)}$, (25)
which becomes the particle filter’s estimation state for object $i$ at time $t$. If $M<5$ (insufficient history) we skip resampling and simply set $\hat{v}_{i,t}=v_{i,t}$.

### v3

@@ Abstract
Augmented Reality (AR) enriches human perception by overlaying virtual elements onto the physical world. However, this tight coupling between virtual and real content makes AR vulnerable to cognitive attacks: manipulations that distort users’ semantic understanding of the environment. Existing detection methods largely focus on visual inconsistencies at the pixel or image level, offering limited semantic reasoning or interpretability. To address these limitations, we introduce CADAR, a neuro-symbolic framework for cognitive attack detection in AR that integrates neural and symbolic reasoning. CADAR fuses multimodal vision–language representations from pre-trained models into a perception graph that captures objects, relations, and temporal contextual salience. Building on this structure, a particle-filter-based statistical reasoning module infers anomalies in semantic dynamics to reveal cognitive attacks. This combination provides both the adaptability of modern vision–language models and the interpretability of probabilistic symbolic reasoning. Preliminary experiments on an AR cognitive-attack dataset demonstrate consistent advantages over existing approaches, highlighting the potential of neuro-symbolic methods for robust and interpretable AR security.
@@ 1 Introduction
Augmented Reality (AR) and Mixed Reality (MR) technologies increasingly overlay virtual content onto the physical world, reshaping how users perceive and interact with their surroundings [zheng2023review]. This tight integration creates new vulnerabilities, as users tend to trust overlaid visual and spatial cues. Adversaries can exploit this trust to launch cognitive attacks, where malicious virtual content is injected or altered to mislead, distract, or manipulate users without their awareness [10924575, cheng2023exploring, teymourian2025sok, xiu2025detectingvisualinformationmanipulation]. Unlike traditional cybersecurity threats that target system integrity, cognitive attacks operate at the perceptual level, subtly manipulating users’ semantic understanding of a scene (examples in Fig.1).
@@
Figure 1: Illustrative AR/MR cognitive attacks. Top row: original scenes; bottom row: manipulated scenes. (a) Text Modification Attack. (b) Visual Modification attack. (c) Obstruction Attack. (d) Injection Attack.
@@
Prior work on detecting anomaly and attacks often employs computer vision techniques to identify visual changes or adversarial distortions [qiu2019review, gu2024anomalygpt, qu2023towards, wang2022objectformer, duan2022confusing, li2024survey, acharya2022detecting]. However, these methods rely on pixel/image processing, lacking the ability to model the human perception process that not only involves prior knowledge and contextual information, but also selectively focuses on aspects of the visual field for semantical interpretation. As a result, these methods may struggle to capture smart cognitive attacks manipulating only a small number of pixels and also may suffer from false positives due to benign visual perturbations that do not affect users’ semantic interpretation [li2024survey]. More recent studies propose pre-trained vision–language models (VLMs) [xiu2025viddar, roy2025zero] can achieve semantic-level detection by processing visual features and incorporating human-like knowledge. But the reasoning of pre-trained VLMs is a black-box approach offering limited interpretability. It is difficult to leverage domain-specific knowledge or model scene/context evolutions or temporal correlations in cognitive attack detection.
@@
In this paper, we consider cognitive attacks that aim to manipulate users’ semantic perception in AR/MR. We propose CADAR (Cognitive Attack Detection in Augmented Reality), a neuro-symbolic approach encompassing two key components: (i) a symbolic perception-graph model, inspired by knowledge and scene graphs [hogan2021knowledge, chang2021comprehensive, yun2025missiongnn, shah2022lmnavrobotic]. It extracted using pre-trained neural VLM models to represent scene/context evolution at the semantic-level, and (ii) a particle-filter model, which is commonly used in non-linear and stochastic systems. It applies a sequential Monte Carlo method to the perception graph for reasoning and detecting cognitive attacks, stabilizes inherently noisy or inconsistent VLM outputs, and mitigates errors or missed detections through probabilistic fusion. It also offers granular, flexible temporal detection, using reference sequences, frame-wise inference localizes attacks per frame, to produce a variable number of labels when needed. Thus, the CADAR approach inherits the adaptability of the pre-trained VLM (neural part), as well as the interpretability and reasoning rigor of particle filtering (symbolic part). It semantically interprets and models the sequential scene changes for robust cognitive attack detection in AR/MR, overcoming existing work’s limitations.
We summarize our contributions as follows:
@@
  • Neuro-symbolic framework. We introduce a hybrid architecture that converts VLM outputs into symbolic perception graphs for explicit and interpretable reasoning in AR cognitive attack detection.
  • Perception graph schema. We design a spatio-temporal graph representation that integrates multimodal information from image/text encoders and detection modules to capture scene evolution.
  • Particle-filter reasoning. We apply particle filtering to symbolic perception graphs, improving robustness to noisy VLM outputs and enabling principled statistical anomaly detection.
  • CADAR-50K dataset. We build the first public AR cognitive attack dataset and demonstrate that CADAR achieves 74.6% accuracy, significantly outperforming all baseline methods.
@@
The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 presents our method; Section 4 reports experimental results; and Section 5 concludes the paper.
@@ 2 Related Work
One well-studied category of visual attacks is adversarial attacks, which introduce imperceptible yet malicious perturbations to input data to mislead machine learning models [zhang2024adversarial, mumcu2024multimodal]. In contrast, cognitive attacks target perception-level semantics and aim to alter a human user’s understanding of a scene [cheng2023exploring, xiu2025detectingvisualinformationmanipulation]. Augmented Reality (AR) systems, which overlay virtual content onto the physical world to enhance perception and decision-making, are particularly vulnerable to such attacks. However, research on cognitive attacks in AR remains limited. Existing methods often rely on access to the original, unaltered scene for comparison or employ pre-trained VLMs with limited interpretability.
@@
Neuro-symbolic approaches couple deep neural networks with symbolic reasoning–i.e., reasoning over explicitly represented knowledge in formal languages, to capitalize on the complementary strengths of subsymbolic pattern recognition and symbolic compositionality [bhuyan2024neuro, nawaz2025review]. Within knowledge-graph research, most neuro-symbolic work addresses link prediction or graph completion, where the hybrid models offer interpretable rules, guided learning, and more transparent decision pathways [delong2024neurosymbolic]. Yet, like standard supervised techniques, many of these solutions still demand sizable labeled datasets. In contrast, our framework introduces a neuro-symbolic reasoning module that combines explicitly encoded rules with statistics, achieving high interpretability while remaining data-efficient.
@@
Particle filter is a sequential Monte Carlo algorithms for recursive Bayesian estimation [chen2003bayesian], it propagates a set of weighted samples to approximate the posterior state distribution, outperforming classical Kalman filters in strongly non-linear or non-Gaussian settings [kunsch2013particle, 5985520]. They are widely used in surveillance, robotics, and navigation, leveraging visual, geometric, and motion cues such as color, texture, and shape [awal2023particle]. In CADAR, we model graph nodes and edges as stochastic particles, enabling robust reasoning under noisy or inconsistent VLM outputs.
@@ 3 Methodology
Figure 2: A system overview of CADAR. Sequential video frames are first transformed into spatial-temporal symbolic perception graphs. The attack-detection module then analyzes these graphs to identify, classify, and localize adversarial attacks, such as the visual modification and injection attacks shown in the examples.
@@ 3.1 Cognitive Attacks
We investigate four representative cognitive attack types that deliberately compromise the semantic integrity of AR scenes (Fig.1). The text-modification attack manipulates on-scene text—such as changing “No Parking” to “Free Parking”—to reverse intended meanings or inject false instructions, thereby deceiving both human users and perception systems. The visual-modification attack distorts object appearance or placement—for example, turning a green light to red or relocating a stop sign—leading to recognition errors and misinformed user decisions. The obstruction attack targets critical information by occluding or deleting essential cues like exit signs, disrupting safety awareness and expected graph relations. Finally, the injection attack introduces fictitious elements, such as fake hazard symbols or virtual labels, that embed misleading cues, divert user attention, and corrupt subsequent reasoning processes.
@@ 3.2 Perception Graph Model
Figure 3: Symbolic perception graph generation: given a video frame and its contextual description, VLMs generate the corresponding perception graph at time step t.
@@
A perception graph consists of nodes and edges. Objects are represented as nodes, and object relationships are represented as edges. Given the context information and prompt (shown in the Fig.3), the VLM would extract the related objects with their relationships in the scene. At timestep $t$, the Perception Graph $G_{t}$ is generated as follows.
3.2.1 Graph Construction.
@@
Thus, the GraphSet $\hat{G}_{t}$ is the union of all spatial and temporal connections that stitch object instances across time steps 1:t1{:}t. Each node still carries its original timestamp to preserve the temporal ordering.
@@
$\hat{G}_{t}=\bigl(\mathcal{V}_{1:t},\,\mathcal{E}_{1:t}\bigr),\hskip 2.84526pt\mathcal{V}_{1:t}=\!\!\bigcup_{i\in\mathcal{I}_{t}}\!\!\mathcal{N}_{i,t},\hskip 2.84526ptE_{1:t}=\!\!\bigcup_{i,j\in\mathcal{I}_{t}}\!\!\mathcal{E}_{ij,t}.$ (7)
@@ 3.3 Particle Filter Model
Our particle filter model tracks graph evolution and analyzes adversarial perturbations. It proceeds in three stages: prediction and matching, attack detection, and state estimation (Fig. 4). Each entity (node or edge) is treated as a particle At each timestep, particles match the incoming observations with prior particles. The attack detection algorithm applies statistical tests to newly matched particles and discards those marked as attacked. Finally, weighting, resampling, and estimation produce an updated, temporally consistent graph state, preserving persistent NodeSet/EdgeSet identities and the relational structure over time.
@@
$s_{j,i,m}\;=\;\frac{1}{|\mathcal{N}_{i,t-1}|}\sum_{v\in\mathcal{N}_{i,t-1}}\operatorname{sim}_{m}(v_{j}^{t},v),\quad m\in\{n,d,f\}.$ (9)
@@
Finally, the assignment rule based on measuring the remaining candidate distance:
$\mu_{t}(v_{j,t})=\begin{cases}\displaystyle\arg\min_{i\in\mathcal{C}_{j}}\sum_{m\in\{n,d,f\}}\bigl|s_{j,i,m}-\mu_{i,m}^{\,t-1}\bigr|,&\mathcal{C}_{j}\neq\varnothing,\\[6.0pt] \text{new\_id},&\mathcal{C}_{j}=\varnothing.\end{cases}$ (11)
If $\mu_{t}(v_{j}^{t})$ returns an existing ID, append $v_{j}^{t}$ to that NodeSet and update $A_{i}^{t}$; otherwise, spawn a new NodeSet. Confirmed edges $e_{ab}^{t}$ are then inserted into $E_{\mu_{t}(a)\,\mu_{t}(b),t}$ as before. After matching, the NodeSet goes into the Attack detection process.
3.3.2 Attack Detection.
@@
where $k_{m}$ is chosen such that $\geq\!99\%$ variance is retained. the resulting embeddings are modeled as Gaussian distributions:
$\mathbf{z}_{m}(v)=\mathbf{W}_{i,m}^{\,t-1\,\top}\mathbf{m}(v)\;\sim\;\mathcal{N}\!\bigl(\mathbf{c}_{i,m}^{\,t-1},\,\boldsymbol{\Sigma}_{i,m}^{\,t-1}\bigr)$ (13)
@@
$\mathbf{c}_{i,m}^{\,t-1}=\frac{1}{|\tilde{\mathcal{N}}_{i,t-1}|}\sum_{v\in\tilde{\mathcal{N}}_{i,t-1}}\mathbf{z}_{m}(v),\hskip 2.84526pt\boldsymbol{\Sigma}_{i,m}^{\,t-1}=\mathrm{Cov}\!\bigl\{\mathbf{z}_{m}(v)\bigr\}.$ (14)
@@
For the structural attacks, specifically object obstruction (Type 3) and object injection (Type 4), the detector relies on lightweight temporal heuristics rather than statistical distance. After each matching step, the algorithm keeps a two-frame sliding window of NodeSet activity. If an object that previously carried a high task-importance score ($\pi\geq\pi_{\text{high}}$) fails to obtain a match in both the current and the immediately preceding frame, the system labels the disappearance an obstruction attack. Conversely, any NodeSet that first appears within the last two frames and whose reasonability score falls below a predefined threshold ($\rho\leq\rho_{\text{low}}$) is treated as an injection attack.
@@
So $w_{j}$ is the normalized weight of the particle, based on its distance to the new particle. Similarly, for embedding attributes, modality $m\in\{n,d,f\}$ (name, description, image) let ${m}^{\star}={m}(v_{i}^{t})$ and ${m}^{(j)}={m}(v^{(j)})$. Weight is defined as
$sim_{j}=\frac{{m}^{(j)}\!\cdot\!{m}^{\star}}{\|{m}^{(j)}\|\,\|{m}^{\star}\|},\qquad w_{j}=\frac{\exp(\lambda_{m}sim_{j})}{\sum_{k=1}^{M}\exp(\lambda_{m}sim_{k})},$ (19)
with a scale factor $\lambda_{m}>0$.
Then use the weight to resample $M$ indices $s_{1},\dots,s_{M}\sim\operatorname{Multinomial}(w_{1:M})$ yields estimated scalar particle and embedding particle:
$\hat{a}\;=\;\frac{1}{M}\sum_{k=1}^{M}a^{(s_{k})},\quad\hat{{m}}\;=\;\frac{\sum_{k=1}^{M}{m}^{(s_{k})}}{\Bigl\|\sum_{k=1}^{M}{m}^{(s_{k})}\Bigr\|},$ (20)
@@
[…]
which becomes the particle filter’s estimation state for object $i$ at time $t$. If $M<5$ (insufficient history), we skip resampling and simply set $\hat{v}_{i,t}=v_{i,t}$.